OpenStudy (jettddl):

Juliana's cell phone bill includes a fixed montly charge of $59/month plus a data charge of $6/GB of data. ?W?hat percent of Julian's cell phone bill for August wasf from the data charge? Round your answer to the nearest tenth of a percent.

OpenStudy (photon336):

is there anymore information to this?

OpenStudy (jettddl):

Nope

OpenStudy (jettddl):

Oh wait actucally Julianna used 4G

OpenStudy (jettddl):

*actually

OpenStudy (photon336):

Well for the bill \[59t = P(t) \] and for the data \[6x = P(x)\] now for august right we know that each month he needs to pay 59 dollars and he also needs to pay 6 dollars per GB of data. my other issue is that i'm asking if there was any more information to the question because how much data did he use? \[%~spent~data~charge = \frac{ bill~data~charge }{ total~bill }*100\]

OpenStudy (jettddl):

4

OpenStudy (photon336):

so he used four GB of data right? so, based on the equations I wrote how much would that be in total for the bill on Data?

OpenStudy (jettddl):

yes. he did but i stil don't really understand

OpenStudy (photon336):

alright so we know that his bill is going to be 59 dollars a month right?

OpenStudy (jettddl):

yup

OpenStudy (photon336):

and that he used 4GB of data

OpenStudy (photon336):

And that they charge $6 per GB that he uses. so how much would this bill be if he used 4 GB of data?

OpenStudy (jettddl):

24

OpenStudy (photon336):

okay now in that month august his bill is 56 dollars. now to find the % he spent on data in that month what we need to do is this: \[\frac{ number~spent~on~data }{ total~bill~data+bill }*100\]

OpenStudy (photon336):

Everything clear?

OpenStudy (jettddl):

i got 30 as an answer

OpenStudy (jettddl):

is it correct @Photon336

OpenStudy (photon336):

30%, yeah